# https://sarahweiii.github.io/meshlrm/ MeshLRM: Large Reconstruction Model for High-Quality Meshes

### v1

@@ 4.3 Comparison Against Baselines
Comparisons with feed-forward methods. To the best of our knowledge, the LRM model proposed in Instant3D [30] is the only previous model that can also achieve 3D object reconstruction from four sparse views. As reported by Instant3D, previous cost-volume-based feed-forward methods (like SparseNeus[38]) cannot handle this challenging case. Therefore, we focus on comparing with the LRM model in Instant3D (labeled as In3D-LRM) on the GSO dataset, showing quantitative results with detailed training settings and timings in Tab. 5. In the table, we show our results from both stages and compare them with both In3D-LRM and In3D-LRM + MC; for In3D-LRM + MC, Marching Cubes is directly applied to extract meshes from their NeRF reconstruction.
As shown in Tab. 5, our model, across both stages, achieves superior quality compared to Instant3D. As expected, In3D-LRM + MC leads to much worse rendering quality than In3D-LRM, due to the quality drop caused by MC. Thanks to our effective 2nd-Stage DiffMC-based training, MeshLRM can reconstruct significantly better meshes in terms of both rendering quality (with 4.2dB PSNR and 0.05 SSIM improvements) and geometry quality (lowering the CD by 0.72). Our mesh rendering is even better than their volume rendering (In3D-LRM), notably increasing their PSNR by 1.4dB and SSIM by 0.03, while achieving substantially faster rendering speed because of outputting meshes. Our superior quality is also reflected by the qualitative comparisons in Fig. 4, where our mesh renderings reproduce more texture and geometric details than the baselines. From Tab. 5, we can also see that our high reconstruction quality is even achieved with a smaller model size and substantially less compute, less than half of the total GPUxDay compute cost required for Instant3D-LRM. This is enabled by our simplified LRM architecture designs and efficient low-res-to-high-res 1st-stage training strategy, leading to significantly faster training speed and convergence. Besides, our model also leads to a faster inference speed, enabling high-quality mesh reconstruction within 1 second. Overall, our MeshLRM leads to state-of-the-art sparse-view mesh reconstruction quality and achieves high efficiency in parameter utilization, and training, inference, and rendering speed. Meanwhile, our MeshLRM (NeRF) has good performance as well and is a state-of-the-art LRM model for sparse-view NeRF reconstruction.
@@
Comparisons with per-scene optimization methods. We also compare our MeshLRM with recent per-scene optimization methods that are designed for sparse-view NeRF reconstruction, including ZeroRF [54] and FreeNeRF [74]. Because of using per-scene optimization, these methods are much slower, requiring tens of minutes and even several hours to reconstruct one single scene. As a result, it is not practical to evaluate them on the GSO dataset, comprising more than 1000 objects. Therefore, we conduct this experiment on the NeRF-Synthetic and OpenIllumination datasets, following the settings used in ZeroRF [54] with four input views. Tab. 6 shows the quantitative novel view synthesis results, comparing our mesh rendering quality with their volume rendering quality. Since ZeroRF also offers a post-processing step to improve their mesh reconstruction from Marching Cubes, we also compare our mesh geometry with theirs using the CD metric on the NeRF-Synthetic dataset (where GT meshes are accessible).
@@
As shown in Tab. 6, our NeRF-Synthetic results outperform FreeNeRF significantly; we also achieve comparable rendering quality and superior geometry quality to ZeroRF. On the challenging OpenIllumination real dataset, our approach outperforms both FreeNeRF and ZeroRF by a large margin, consistently for all three metrics. The OpenIllumination dataset, consisting of real captured images, demonstrates our model’s capability to generalize to real captures for practical 3D reconstruction applications, despite being trained only on rendered images. In addition to our superior (or comparable) quality, our feed-forward mesh reconstruction approach is significantly faster than these optimization-based NeRF methods, in terms of both reconstruction (1000 to 10000 speed up) and rendering.
5 Applications
Our approach achieves efficient and high-quality mesh reconstruction in less than one second, thus naturally enabling many 3D generation applications, when combined with multi-view image generation techniques.
Text-to-3D Generation We apply the multi-view diffusion models from Instant3D [30] to generate 4-view images from text inputs, followed by our Mesh-LRM to achieve text-to-3D generation. Our results and comparison with the original Instant3D pipeline with In3D-LRM (that is already quantitatively compared in Tab. 5 in terms of a reconstruction task) are shown in Fig. 5. Our approach leads to better mesh quality and fewer rendering artifacts.
Image-to-3D Generation We apply Zero123++ [53] v1.2 to generate multi-view images from a single image input, followed by our MeshLRM to achieve image-to-3D generation. Our results and comparison with the One2345++ [34], which uses a diffusion-based reconstruction model, are shown in Fig. 6. Our results are of much higher quality, containing sharper textures and more faithful geometric details than One2345++.
6 Conclusion
In this paper, we have presented MeshLRM, a novel Large Reconstruction Model that can directly output high-quality meshes. We achieved it by applying the Differentiable Marching Cubes (DiffMC) method and differentiable rasterization to fine-tune a pre-trained NeRF-based LRM trained with volume rendering. As DiffMC requires backbone efficiency, we brought multiple improvements (tiny shared MLP and simplified image tokenization) to the LRM architecture, facilitating both NeRF and mesh training. We also found that a low-to-high-res training strategy can significantly accelerate the training of the NeRF-based model. Compared with existing methods, our method has both quality increase and speed improvement and is the only one that can output high-quality meshes. In addition, we showed that our method can be directly applied to applications like text-to-3D and image-to-3D generation. As meshes are the most widely accepted format for 3D assets in the industry, we believe our method takes a step towards automating 3D asset creation and potentially opens up new possibilities with novel 3D workflows.
Acknowledgement
Thanks to Nathan Carr for his support on this project and to Linghao Chen for helping to run the One2345++ baseline.
Appendix 0.A Effectiveness of surface fine-tuning (Stage 2).
@@ Acknowledgement
We have discussed the effectiveness of our second-stage fine-tuning in Sec. 4.2 with quantitative results shown in Tab. 2 in the main paper. We now show examples of qualitative results in Fig. 7, comparing our final meshes (MeshLRM) with the meshes (MeshLRM (NeRF) + MC) generated by directly applying Marching Cubes on the 1st-stage model’s NeRF results. As shown in the figure, the MeshLRM (NeRF) + MC baseline can lead to severe artifacts with non-smooth surfaces and even holes, whereas our final model with the surface rendering fine-tuning can effectively address these issues and produce high-quality meshes and realistic renderings. These qualitative results demonstrate the large visual improvements achieved by our final model for mesh reconstruction, reflecting the big quantitative improvements shown in the paper.
Appendix 0.B ABO and OpenIllumination datasets
@@
We compare our proposed method with In3D-LRM [30] on 1000 randomly sampled examples in the ABO [12] dataset with quantitative results shown in Tab. 7. In particular, ABO is a highly challenging synthetic dataset that contains many objects with complex glossy materials. This is a big challenge for our model, as well as In3D-LRM’s model, since both assume Lambertian appearance. In this case, the NeRF models are often better than the mesh models, since NeRF rendering can possibly fake glossy effects by using colors from multiple points along each pixel ray while mesh rendering cannot. Nonetheless, as shown in the table, both our NeRF and final mesh model can outperform In3D-LRM’s NeRF and mesh variants respectively. Especially, thanks to our effective MC-based finetuning, our mesh rendering quality surpasses that of In3D-LRM + MC by a large margin (e.g. 3.92dB in PSNR and 0.48 in SSIM).
On the other hand, we also compare with In3D-LRM on the full OpenIllumination [32] dataset in Tab. 7. To clarify, this experiment setting is different from the setting used in Tab. 6 in the main paper, where we follow ZeroRF [54] to test only 8 objects with combined masks in OpenIllumination; here, we evaluate all models on the full dataset with 100 objects with object masks and we also take square crops that tightly bounding the objects from the rendered/GT images to compute rendering metrics, avoiding large background regions. As shown in the table, we observe that In3D-LRM cannot generalize to handle this challenging real dataset, leading to very low PSNRs, despite being trained on the same training data as ours. In contrast, our model can still work well on this out-of-domain dataset with high rendering quality.
Appendix 0.C Additional Implementation Details
The transformer size follows the transformer-large config [15]. It has layers and a model width of . Each attention layer has a attention head and each head has a dim of . The intermediate dimension of the MLP layer is . We use GeLU activation inside the MLP and use Pre-LN architecture. The layer normalization is also applied after the input image tokenization, and also before the triplane’s unpachifying operator.
@@ Appendix 0.C Additional Implementation Details
Appendix 0.D Limitations
Since our model employs surface rendering and assumes only Lambertian appearance without performing any inverse rendering, it is not sufficiently robust when the input images contain complex materials, e.g. some examples in the ABO dataset. The generated mesh may use white color to fit the specular areas. We believe that incorporating inverse rendering into the current pipeline could address this issue. However, this requires sufficient training data with accurate material annotations, which we leave for future exploration. On the other hand, while handling highly sparse input views, our model requires input camera poses. Although poses can be readily obtained from text- or image-to-multi-view models [30, 53] for 3D generation tasks, calibrating sparse-view poses for real captures can be a big challenge. In the future, it will be an interesting direction to explore combining our approach with recent pose estimation techniques [64, 66] for joint mesh reconstruction and camera calibration.

### v2

@@ Appendix B Comparison with per-scene optimization methods
We compare our feed-forward sparse-view reconstruction method with FreeNeRF (Yang et al., 2023) and ZeroRF (Shi et al., 2023b) on NeRF-Synthetic (Mildenhall et al., 2020) and OpenIllumination datasets (Liu et al., 2024a).
@@
Due to their use of per-scene optimization, these methods are much slower, requiring tens of minutes and up to several hours to reconstruct a single scene. As a result, it is not practical to evaluate them on the GSO dataset, comprising more than 1000 objects. Therefore, we conduct this experiment on the NeRF-Synthetic and OpenIllumination datasets, following the settings used in ZeroRF (Shi et al., 2023b) with four input views. Tab. 7 shows the quantitative novel view synthesis results, comparing our mesh rendering quality with their volume rendering quality. Since ZeroRF uses a post-processing step to improve their mesh reconstruction from Marching Cubes, we compare our mesh geometry with theirs using the CD metric on the NeRF-Synthetic dataset (where GT meshes are accessible).
@@
As shown in Tab. 7, our NeRF-Synthetic results significantly outperform FreeNeRF; we also achieve comparable rendering quality and higher geometry quality to ZeroRF in a fraction of the execution time. On the challenging OpenIllumination real dataset, our approach outperforms both FreeNeRF and ZeroRF by a large margin, consistently for all three metrics. The OpenIllumination dataset, consisting of real captured images, demonstrates our model’s capability to generalize to real captures for practical 3D reconstruction applications, despite being trained only on rendered images. In addition to our superior quality, our feed-forward mesh reconstruction approach is significantly faster than these optimization-based NeRF methods, in terms of both reconstruction (1000 to 10000 speed up) and rendering.
Appendix C Application visualizations
Our feed-forward reconstruction model can be used for 3D generation from text or single image prompts. We apply the multi-view diffusion models from Instant3D (Li et al., 2024) for text-to-multi-view generation and Zero123++ (Shi et al., 2023a) for image-to-multi-view generation. A visual comparison of our method with other 3D generation methods is shown in Fig. 7 and Fig. 6.
Appendix D Inference on various numbers of views
Although our model is trained with only 4 input views, it is capable of handling varying numbers of views during inference. We evaluate MeshLRM on 2-8 input views using the GSO dataset, as presented in Tab. 8. As the number of input views increases, the model demonstrates consistent performance improvements, while still producing strong results even with just two input views, highlighting the robustness and flexibility of our approach.
Appendix E Effectiveness of removing DINO encoder.
We evaluate our model alongside In3D-LRM Li et al. (2024) (both trained on 512512 resolution images) on 10241024 images, as shown in Fig 8. We observe significant geometry distortion in the In3D-LRM results, while our model still produces reasonable outputs. This demonstrates that our choice of positional encoding, which eliminates the DINO encoder and relies solely on Plücker Rays, greatly improves the model’s robustness to changes in image resolution.
| # views | 2 | 3 | 4 | 6 | 8 |
| PSNR | 24.85 | 26.91 | 27.93 | 29.09 | 29.35 |
| SSIM | 0.897 | 0.916 | 0.925 | 0.935 | 0.938 |
| LPIPS | 0.099 | 0.087 | 0.081 | 0.071 | 0.069 |
Appendix F Effectiveness of surface fine-tuning (Stage 2).
@@ Appendix D Inference on various numbers of views
We discuss the effectiveness of our second-stage fine-tuning in Sec. 4.2 with quantitative results shown in Tab. 2 in the main paper. We now show examples of qualitative results in Fig. 9, comparing our final meshes (MeshLRM) with the meshes (MeshLRM (NeRF) + MC) generated by directly applying Marching Cubes on the 1st-stage model’s NeRF results. As shown in the figure, the MeshLRM (NeRF) + MC baseline leads to severe artifacts with non-smooth surfaces and even holes, while our final model with the surface rendering fine-tuning can effectively address these issues and produce high-quality meshes and realistic renderings. These qualitative results demonstrate the large visual improvements achieved by our final model for mesh reconstruction, reflecting the big quantitative improvements shown in the paper.
Appendix G ABO and OpenIllumination datasets
@@
We compare our proposed method with In3D-LRM (Li et al., 2024) on 1000 randomly sampled examples in the ABO (Collins et al., 2022) dataset with quantitative results shown in Tab. 9. In particular, ABO is a highly challenging synthetic dataset that contains many objects with complex glossy materials. This is a big challenge for our model, as well as In3D-LRM’s model, since both assume Lambertian appearance. In this case, the NeRF models are often better than the mesh models, since NeRF rendering can possibly fake glossy effects by using colors from multiple points along each pixel ray while mesh rendering cannot. Nonetheless, as shown in the table, both our NeRF and final mesh model can outperform In3D-LRM’s NeRF and mesh variants respectively. Especially, thanks to our effective MC-based finetuning, our mesh rendering quality surpasses that of In3D-LRM + MC by a large margin (e.g. 3.92dB in PSNR and 0.48 in SSIM).
We also compare with In3D-LRM on the full OpenIllumination (Liu et al., 2024a) dataset in Tab. 9. This experiment setting is different from the setting used in Tab. 6 in the main paper, where we follow ZeroRF (Shi et al., 2023b) to test only 8 objects with combined masks in OpenIllumination; here, we evaluate all models on the full dataset with 100 objects with object masks and take square crops that tightly bound the objects from the rendered/GT images to compute rendering metrics, avoiding large background regions. As shown in the table, we observe that In3D-LRM cannot generalize to this challenging real dataset, leading to very low PSNRs, despite being trained on the same training data as ours. In contrast, our model still works well on this out-of-domain dataset with high rendering quality.
Appendix H Additional Implementation Details
The transformer size follows the transformer-large config (Devlin et al., 2019). It has layers and a model width of . Each attention layer has a attention head and each head has a dim of . The intermediate dimension of the MLP layer is . We use GeLU activation inside the MLP and use Pre-LN architecture. The layer normalization is also applied after the input image tokenization, and before the triplane’s unpachifying operator.
@@ Appendix H Additional Implementation Details
For the CD metric, we sample points only on the visible surface from the testing views. Since the ground-truth meshes from GSO Downs et al. (2022) and OmniObject3D Wu et al. (2023) are obtained through scanning, their interior structures cannot be considered reliable. For example, we found that the interiors of shoes in GSO are incorrectly represented, with a face filling the interior, which contradicts common sense. To address this, we cast rays from all testing views and sample 100,000 points at the ray-surface intersections for each object, following a strategy similar to that used in Tang et al. (2022).
Appendix I Details for image-to-3D testing setup
For the GSO dataset, the single-view input is the first thumbnail image, while for the OmniObject3D dataset, the input is a rendered image with a randomly selected pose. We select 24 camera poses evenly distributed around the object to capture a full 360-degree view and use BlenderProc to render images at a resolution of 320×320. The object is normalized to fit within a bounding box. The camera positions are defined by azimuth angles of [0, 45, 90, 135, 180, 225, 270, 315] degrees and elevation angles of [60, 90, 120] degrees.
Appendix J Limitations
As our model employs surface rendering and assumes Lambertian appearance without performing inverse rendering, it is not sufficiently robust when the input images contain complex materials, such as the metal objects shown in Fig. 10. The generated mesh may bake the shadow and use white color to fit the specular areas. We believe that incorporating inverse rendering into the current pipeline could address this issue. However, this requires sufficient training data with accurate material annotations, which we leave for future exploration. On the other hand, while handling highly sparse input views, our model requires input camera poses. Although poses can be readily obtained from text- or image-to-multi-view models (Li et al., 2024; Shi et al., 2023a) for 3D generation tasks, calibrating sparse-view poses for real captures is a challenge. In the future, it will be an interesting direction to explore combining our approach with recent pose estimation techniques (Wang et al., 2023a; b) for joint mesh reconstruction and camera calibration.